Differentiate: y=(1-secx)/(tanx)
1
Expert's answer
2010-09-21T09:38:44-0400
y´=((1-secx)/(tanx))´=((1-(1/cosx))/(sinx/cosx))´=((cosx-1)/sinx)´=(-sinx*sinx-cosx*(cosx-1))/sin²x=(-sin²x-cos²x+cosx)/sin²x=(cosx-1)/sin²x

Answer:(cosx-1)/sin²x